OpenStudy (twizttiez):

Which number is irrational. A. 0.14 sorry for not knowing math terms and such but there's a bar above the "14" B. 1/3 C. pi 4 D. pi 6

OpenStudy (twizttiez):

@mathmate

OpenStudy (mathmate):

Do you know what is an irrational number?

OpenStudy (twizttiez):

Isn't it a number that can be put in decimal form or percentage form?

OpenStudy (twizttiez):

can't*

OpenStudy (twizttiez):

Oops

OpenStudy (mathmate):

In mathematics, many of the questions can be answered correctly if we know the definition of the terms. If we do not know what the terms mean, the question will appear difficult. So the first step is to understand the meaning of the terms. From wiki: "a rational number is any number that can be expressed as the quotient or fraction p/q of two integers, p and q, with the denominator q not equal to zero." " an irrational number is any real number that cannot be expressed as a ratio of integers. Irrational numbers cannot be represented as terminating or repeating decimals."

OpenStudy (mathmate):

Are the terms clear to you now?

OpenStudy (twizttiez):

Somewhat

OpenStudy (mathmate):

For example: (not the same as the first option) 0.14 = 14/100 =7/50, so it can be expressed as a fraction. is it rational?

OpenStudy (twizttiez):

Yes

OpenStudy (twizttiez):

Right?

OpenStudy (mathmate):

Good!

OpenStudy (twizttiez):

Okay

OpenStudy (twizttiez):

So a number that can become a Fraction or a decimal correct?

OpenStudy (mathmate):

Now the number 0.13333333333... is called a repeting decimal. It can be converted to a fraction as 2/15

OpenStudy (mathmate):

All repeting decimals can be converted to a fraction.

OpenStudy (twizttiez):

Oooookay

OpenStudy (twizttiez):

So we're looking for a number that can't be turned into a fraction right?

OpenStudy (mathmate):

Right! If you need further help, let me know.

OpenStudy (twizttiez):

Oops

OpenStudy (twizttiez):

Option C and D were supposed to be 4 to the fourth and 6 to the fourth

OpenStudy (mathmate):

I knew there was a typo! So are you all good now?

OpenStudy (twizttiez):

Yes I think it would be D I don't know if I'm correct though

OpenStudy (twizttiez):

@mathmate

OpenStudy (mathmate):

To avoid communication errors, please give me the answer and not the letter.

OpenStudy (twizttiez):

Oh okay sorry ^-^

OpenStudy (twizttiez):

6 to the fourth

OpenStudy (twizttiez):

You said that repeating decimals can be turned into fractions and A and B are both repeating

OpenStudy (twizttiez):

And 4 to the fourth is 2

OpenStudy (twizttiez):

You there?

OpenStudy (twizttiez):

@mathmate

OpenStudy (twizttiez):

Hello?

OpenStudy (twizttiez):

@campbell_st

OpenStudy (campbell_st):

well pi is irrational.... so dividing it by anything results in an irrational number.

OpenStudy (twizttiez):

Oh sorry

OpenStudy (twizttiez):

I mistyped earlier

OpenStudy (twizttiez):

4 and 6 are supposed to be put to the fourth

OpenStudy (campbell_st):

that's ok... I think there are 2 irrational numbers in your answer choices... they both involve pi...

OpenStudy (twizttiez):

Yeah as I said C and D aren't supposed to have Pi

OpenStudy (twizttiez):

They're supposed to be put to the fourth

OpenStudy (campbell_st):

can you use the draw function to write them

OpenStudy (twizttiez):

\[\sqrt{4} \sqrt{6}\]

OpenStudy (campbell_st):

I'll let @mathmate continue to answer this

OpenStudy (mathmate):

@TwiztTiez Remember that all integers are rational, because 5 is the same as 5/1, so it can be represented by a fraction. So the 4 choices are: 1.44444444..... (or 1.4 with a rep. dec. bar) 1/3 4^6 6^6 ? That's why I prefer dealing with the actual numbers, not the choice letters.

OpenStudy (mathmate):

or rather 1.44444444..... (or 1.4 with a rep. dec. bar) 1/3 4^6 \(\sqrt4 \sqrt6\) which is the same as \(\sqrt{24}\)

OpenStudy (mathmate):

Which one of these can you NOT represent by a fraction?

OpenStudy (twizttiez):

I didn't mean for it to look like they were supposed to multiplied sorry

OpenStudy (mathmate):

That's ok, the multiplication is understood. I just used the law of exponents to simplify it.

OpenStudy (twizttiez):

Well \[\sqrt{4} =2\]

OpenStudy (mathmate):

For your information 1.4444444... equals 13/9.

OpenStudy (mathmate):

Very good start, continue! So \(\sqrt{4}\sqrt6=2\sqrt{6}\) can you go further?

OpenStudy (twizttiez):

Yup yup

OpenStudy (mathmate):

I'll put it this way, can you represent \(\sqrt6\) by a fraction?

OpenStudy (twizttiez):

4.8989794

OpenStudy (mathmate):

...mathematically exactly?

OpenStudy (twizttiez):

\[2\sqrt{6}=4.8989794...\]

OpenStudy (mathmate):

2.449489742783178.... and this is not exact.

OpenStudy (mathmate):

4.898979485566356... and this is not exact.

OpenStudy (twizttiez):

Well how can I make it exact? Is there a way or am I just being oblivious?

OpenStudy (twizttiez):

And I don't mean that rudely

OpenStudy (twizttiez):

Sorry if it came off that way

OpenStudy (mathmate):

Well, there is no way to represent a square root in decimal form unless it is a perfect square. What does that mean (in the context of this question)?

OpenStudy (mathmate):

4.89897948556635619639456814941178278393189496131334025686539... and this is still not exact!

OpenStudy (twizttiez):

Oh okay so

OpenStudy (twizttiez):

\[\sqrt{4}=2\]

OpenStudy (twizttiez):

That's exact right?

OpenStudy (mathmate):

yes, sqrt(4) is rational.

OpenStudy (twizttiez):

Okay I get it

OpenStudy (mathmate):

but is sqrt(6)?

OpenStudy (mathmate):

It can be proved mathematically that square root of a number (such as 2, 3, 5,6,...) is not rational unless the number is a perfect square (1,4,9,16,25....)

OpenStudy (twizttiez):

2.44948974278

OpenStudy (twizttiez):

So no I don't think so

OpenStudy (mathmate):

You can read more about it (and it is interesting reading for general information) at, for example, Wiki: http://en.wikipedia.org/wiki/Irrational_number or http://www.purplemath.com/modules/numtypes.htm

OpenStudy (twizttiez):

Thanks!

OpenStudy (mathmate):

Correct, since sqrt(6) is not rational, so 2sqrt(6) is also not rational. Does that clear things up a little?

OpenStudy (twizttiez):

It does thanks for the help!

OpenStudy (mathmate):

Good, you're welcome! :)